.A farmer produces seeds in packets for sale. The probability that a seed selected will grow is 0.8. If 5 of the seeds are grown, what is the probability that?:

i.less than two will grow? 

ii.less more than three will grow?


1
Expert's answer
2021-07-20T09:13:43-0400

Binomial (5,0.8)

"P(x)=\binom{n}{x}P^x(1-P)^{n-x}, x=0,1,...,5."


I "P(<2)"

"P(x<2)=P(x=0)+P(x=1)"

"=\binom{5}{0}(1-0.8)^{5}+\binom{5}{1}0.8(1-0.8)^{4}"

"=0.00032+0.0064"

"=0.672"


Ii. P(<3)

"P(x<3)=P(x=0)+P(x=1)+P(x=2)"

From part one, we already have "P(x=1)+P(x=0)=0.00672"

Thus,

"P(x<3)=0.00672+P(x=2)"

"=0.00672+\binom{5}{2}0.8^2(1-0.8)^{3}"

"=0.0672+0.0512"

"=0.05792"